Sol Summary Report – March 7th

Crew 275 Sol Summary Report 07Mar2023

Sol: 23

Summary Title: Keeping the (space)ship afloat!

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

With three crewmembers leaving to the hospital early in the morning, only four crewmembers remained at the Hab today. Since we knew that things will be more difficult in this context, especially with the presence of journalists onsite, we planned mainly routine activities.

Due to the late arrival of the journalist, we had to slightly delay the EVA dedicated to changing the batteries of the atmospheric instruments and retrieving their data. As planned, we also took some time to go to North Ridge and shoot nice videos with the journalists, including drone footage. The afternoon was then mainly dedicated to showcasing our different experiments and touring the different buildings of the campus with the journalists.

With the temperature in the GreenHab close to 40°C/100°F, we noticed that the leaves of the tomato crops were turning yellow. We are also very sad to announce that one of our fishes has passed out for an unknown reason. The water tests seem to indicate that this event was not linked to the quality of the water. The GreenHab officer will keep an eye on this in the next few days.

Some news also from our injured crewmember. After another X-ray and an appointment with a surgeon in the end of the morning, they have directly proceeded with surgery in the afternoon. It was a success, and they feel well. Everyone feels relieved to know that the three missing crewmembers will be back to the Hab in a few hours!

Look Ahead Plan:

With the three missing crewmembers back to the Hab in a few hours, we will be able to go back to our nominal planning. No EVA is planned for tomorrow, but we would like to perform an exploration EVA on Thursday (Sol 25) and another EVA dedicated to the disassembling of the atmospheric instruments on Friday (Sol 26).

Anomalies in work: None.

Weather: Sunny, low -3°C/26°F, high 11°C/52°F.

Crew Physical Status: One crewmember injured (broken collarbone / clavicula). Together with the Health & Safety Officer and the Crew Scientist, they left for the hospital in the morning to see a surgeon at the Sevier Valley hospital. The surgery was successfully conducted in the afternoon, and they are about to come back to the MDRS.

EVA: EVA # 19 was successfully conducted, after requesting a delay due to the late arrival of the journalist team from France Televisions. The batteries of the atmospheric instruments have been changed and the data retrieved. Then the EVA crew went to North Ridge to shoot some nice videos with the journalists.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site